OpenStudy (anonymous):

Please answer this: The sum of two consecutive positive odd integers is between 7 and 17. Which pair of integers does not solve this problem?

OpenStudy (anonymous):

3 and 5 7 and 9 5 and 7 1 and 5

OpenStudy (anonymous):

add them A) 3 + 5 = 8 between 7 and 17 B) 7+9 = 16 between 7 and 17 C) 5+7 = 12 between 7 and 17 D) 1+ 5 = 6 less than 7

OpenStudy (anonymous):

thanks!

OpenStudy (goformit100):

7 and 9